# InvestESG: A multi-agent reinforcement learning benchmark for studying climate investment as a social dilemma

### v2

@@ 4.1 The trade-off for the agents frames climate change as a social dilemma
Companies’ trade-off between short-term private costs and long-term collective benefits create a social dilemma within the environment. To illustrate this, we use empirical Schelling diagrams (Hughes et al., 2018), which plot the focal-company payoffs from following either a cooperative or defecting policy, depending on the number of other cooperating company agents in a 5-company, 3-investor environment.
@@
Figure 3: Schelling diagrams demonstrating that the environment constitutes a social dilemma. The graphs compare payoffs between cooperation (mitigation, blue lines) and defection (no mitigation, red lines) for a focal company, given varying number of other cooperating companies. Yellow lines represent the average payoff across all companies when the focal company defects. Subfigure (a) illustrates the selfish scenario, where all three investors consistently prioritize financial returns ($\alpha^{\mathcal{I}_{j}}=0,\text{ for }j=1,2,3$). Here, defection always yields higher payoffs for the focal company than cooperation, leading all companies to defect. However, widespread defection results in lower overall profits, as the average payoff (yellow) increases with greater cooperation. Subfigure (b) and (c) correspond to two and three infinitely ESG-conscious investors ($\alpha^{\mathcal{I}_{j}}\approx\infty$), respectively. In (b), cooperation yields higher payoffs than defection for the focal company when few others cooperate. In (c), cooperation outperforms defection in all cases. Therefore, (b-c) demonstrate how investor behavior can transform the environment, eliminating the social dilemma by aligning corporate incentives with mitigation. Subfigures (d) and (e) build on (c) with three ESG-conscious investors. Subfigure (d) introduces resilience spending, while (e) adds greenwashing. The latter reintroduces a social dilemma, where corporations again avoid mitigation.
@@
In our case, cooperation represents altruistic investment in mitigating emissions. We simulate a cooperative policy as one which consistently invests 0.5% [^3] Since company actions are continuous, we selected 0.5% as a benchmark for illustrating the social dilemma structure in the Schelling diagram. This aligns with Patagonia’s cooperative approach, where $\sim$0.3of its valuation is allocated annually to climate initiatives. The diagram’s shape remains consistent across other tested values. of capital in mitigation and 0 in greenwashing or resilience building. In Figure 3(a) to 3(c), a defector policy takes zero action in all of mitigation, greenwashing, and resilience. As shown in Figure 3(a), when investors are solely profit-driven, the payoff for following a defector policy is always higher than the cooperator payoff. However, if all agents fail to cooperate, overall payoffs will be extremely low, showing that the environment represents a social dilemma. Figures 3(b) and 3(c) illustrate how the payoffs for cooperation and defection evolve as the proportion of ESG-conscious investors with positive ESG preferences increases. Notably, Figure 3(b) suggests the possibility of a bifurcated equilibrium when investor preferences are mixed. Specifically, climate-friendly companies may attract ESG-conscious investors, while other companies free ride and attract profit-motivated investors. Figure 3(c) shows that when all investors prioritize ESG-friendly firms, cooperation dominates defection in all cases. Therefore, with enough ESG-conscious investors the environment is no longer a social dilemma, and companies are actually greedily motivated to invest in mitigation to attract investors.
@@ 4.2 Multi-agent Reinforcement Learning Baselines
We are motivated to simulate companies and investors as independent, selfishly motivated agents that specialize in maximizing their own expected reward. We employ a state-of-the-art MARL algorithm based on Independent PPO (IPPO) (De Witt et al., 2020; Yu et al., 2022). so that each agent has its own policy parameters, and agents do not share parameters among themselves. At the back-end, InvestESG is implemented in both PyTorch (Paszke et al., 2017) and JAX (Bradbury et al., 2018), capable of simulating the investment behaviors in a massively parallel sense by vectorization on both CPUs and GPUs.
By default, 5 companies are modeled with an initial capital of \$10 trillion each, alongside 3 investors, each starting with \$16 trillion. This creates a total beginning market wealth of \$98 trillion, which is comparable to the global stock market cap (WFE Statistics, 2023). We conduct the experiments in such a 5-company, 3-investor environment by default unless specified otherwise. All the following results are averaged result based on 3 runs of different random seeds, with the standard error plotted in the shaded region [^5] In simulating climate events as described in Equation equation 2, we use a fixed random seed across training episodes. This approach ensures that, while event occurrence varies with episode-specific climate risks, the underlying randomness remains deterministic. This setup is based on two considerations. First, unlike MARL agents learning from scratch, real-world companies possess a baseline understanding of trade-offs and can reasonably anticipate future risks; a deterministic training environment mirrors this insight. Second, our study aims not to develop agents that solve climate change under all stochastic scenarios, but rather to train agents that grasp the strategic trade-offs faced by companies and investors. A partially deterministic environment makes learning and convergence easier without compromising the study’s structural conclusions. For robustness, the environment setup allows agents to be trained in a fully stochastic environment to ensure generalizability. . Following economic literature that argues more than 10 agents are often needed to observe meaningful group size effects (Arifovic et al., 2023), we provide additional results in Appendix 11 from settings with 10 companies and 10 investors, as well as 25 companies and 25 investors, which are consistent with our findings with the 5-company, 3-investor setting. [^6] The 25-by-25 agent setting aligns with Zhang et al. (2022), which employs 27 agents in a MARL framework to address climate change issues.
@@ 5 Main Results
Figure 4: Climate risk, total market wealth, total mitigation amount and climate event occurrence over the course of training for the 5-company-3-investor case. We compare the status quo scenario with solely profit-driven investors (investors with ESG consciousness level of 0), both with and without the ESG disclosure mandate, to scenarios involving three ESG-conscious investors with ESG consciousness level of $\alpha=0.5$, $\alpha=1$, and $\alpha=10$. These results indicate that merely disclosing ESG scores is insufficient to resolve the social dilemma if investors are not interested in investing in climate-friendly companies. Only when investors have a high level of ESG consciousness—$\alpha=10$, in this case—does the ESG mandate make a difference in increasing the level of mitigation.
Highly ESG-conscious investors are required to effectively incentivize mitigation efforts. To evaluate the effectiveness of ESG disclosure mandates in incentivizing emissions mitigation by companies, we conduct three experiments: (1) Status Quo: All agents are profit-motivated, and no ESG scores are released; (2) Status Quo with Mandate: ESG scores are disclosed, but investors remain profit-driven ($\alpha=0$); (3) Mandate with ESG-Conscious Investors: ESG scores are disclosed, and investors are ESG-conscious ($\alpha>0$). In all scenarios, companies only choose mitigation effort levels, with greenwashing and resilience spending disabled. Figure 4 tracks evaluation metrics over training. Initially, company agents act randomly, leading to frequent mitigation and optimal collective outcomes in the first 200 episodes. As training progresses, agents adopt profit-maximizing strategies, reducing mitigation efforts and achieving sub-optimal collective outcomes characteristic of social dilemmas. Over time, agents gradually realize that, in high-risk scenarios, investing in mitigation becomes marginally beneficial, leading to a slow increase in total market wealth.
The final system climate risks for the Status Quo with Mandate and low-ESG-conscious scenarios ($\alpha=0.5$ and $1$) are similar to those in the Status Quo, indicating that mandatory ESG disclosure alone does not significantly incentivize mitigation when investors prioritize profits. However, when investors have a high level of ESG-consciousness ($\alpha=10$), companies invest significantly in mitigation, reducing climate risk and increasing market wealth. In this setting, a few leading companies attract the majority of ESG-conscious investments, forming a positive feedback loop where better market performance draws further investments (see Appendix Figure 11(c) for a demonstration for such pattern in the 10 companies, 10 investors case).
These findings align with research showing that ESG disclosure mandates encourage climate-friendly efforts driven by societal and stakeholder pressures (Cormier and Magnan, 1999; Gamerschlag et al., 2011; Fama and French, 2007; Friedman and Heinle, 2016; Chen et al., 2018; Ioannou and Serafeim, 2019), and that fund managers are willing to sacrifice financial returns for ESG benefits (Krueger et al., 2021).
(a) Investment matrix for investors with different level of ESG consciousness: 0, 10, 10
@@
Figure 5: Investigating the effects of the level of ESG consciousnessness in the case of 5 companies and 3 investors. ESG consciouness levels are $\alpha^{\mathcal{I}_{0}}=0$, $\alpha^{\mathcal{I}_{1}}=\alpha^{\mathcal{I}_{2}}=10$. In (a) and (b), Company 0 is the leading mitigator out of all the 5 companies. The figure plots the investment distribution for each investor, showing that more climate-conscious investors focus on investing in the more climate-conscious companies, mirroring the market bifurcation results of the Schelling diagrams.
@@
Heterogeneous investor preferences lead to bifurcation in agent strategies. Research shows that investors have varying preferences for ESG efforts and respond in different ways (Amel-Zadeh and Serafeim, 2018). In this scenario, we initialized three investors with different levels of ESG-consciousness, with Investor 0 representing the solely profit-seeking investor with ESG-preference set to $\alpha^{\mathcal{I}_{0}}=0$, and Investor 1 and 2 representing highly ESG-conscious investors with $\alpha^{\mathcal{I}_{1}}=\alpha^{\mathcal{I}_{2}}=10$, which matches the scenario presented in the Schelling diagram in Figure 3(b). Figure 5(a) and 5(b) highlight a divergence in both investor and company behavior. Profit-driven investor 0 distributes investments evenly across companies, while ESG-focused investors (1 and 2) favor climate-conscious firms, such as Company 0, which prioritizes mitigation. Such divergence extends to company strategies: Company 0 learns to attract more ESG investment by focusing on mitigation, while others prioritize financial returns at the expense of some investor interest.
(a) ESG-conscious investors can be distracted by greenwashing.
@@
(b) Mitigation, greenwashing, and climate risk for greenwashing coefficient = 2, 10, 20.
@@
Figure 6: (a) shows ESG-conscious investors can be distracted by greenwashing, and heavily invest in a company that greenwashes. Here we examine a scenario where Company 0 is hard-coded to invest in real mitigation while Company 1 only invests in greenwashing. Investors have ESG consciousness level of 0, 1, 10. Part (b) shows that when both companies and investors are IPPO agents, regardless of greenwashing cost, companies initially explore greenwashing and mitigation equally but quickly abandon greenwashing and invest in mitigation to attract ESG-conscious investors. In these experiments, the final climate risk is similar to the value when greenwashing is not enabled, suggesting greenwashing does not hinder mitigation investment.
The possibility of greenwashing does not significantly undermine mitigation efforts. Building on research that examines the existence and extent of greenwashing (Wu et al., 2020; Marquis et al., 2016; Siano et al., 2017), and the concern that the ESG disclosure policy can backfire with greenwashing (El-Hage, 2021), we explore whether companies adjust their strategies when greenwashing is permitted. Based on the Schelling diagram in Figure 3(d) and Figure 5(b), we predict that investors would be misled by greenwashing, prompting companies to prioritize greenwashing over mitigation due to its lower cost, leading to waste of resources. However, this prediction is not observed when simulating agents with IPPO. Figure 6(b) plots results under varying greenwashing costs, represented by coefficient $\beta$ in Equation 3, where a larger $\beta$ indicates cheaper greenwashing. All investors have an ESG-consciousness level of $\alpha=1$. The results show that regardless of greenwashing cost, companies initially explore greenwashing and mitigation equally but quickly drop greenwashing as the main strategy.
@@
This likely occurs because, in the early episodes, investors have not yet linked ESG scores to their investment strategies, so do not respond to greenwashing efforts with increased investment. Instead, greenwashing presents an immediate cost for companies, which does not pay off with reduced climate risk, so they quickly learn to avoid it. In contrast, even without immediate investor rewards, companies may recognize the long-term benefits of mitigation (which leads to higher collective returns), and be incentivized to continue those efforts at first, even if they later learn to defect. This mirrors reality, where if investors are slow to adjust their strategies, companies may abandon greenwashing early but persist in low levels of mitigation for either its long-term climate benefits or the foresight of regulations and societal pressure that will eventually nudge them towards sustainable operation anyway. This is consistent with some empirical literature stating that despite the emergence of some greenwashing, ESG disclosure mandates overall encourage mitigation (Fiechter et al., 2022).
@@
Providing additional information about climate risk increases mitigation efforts. In this scenario, we want to observe the effect of having additional climate-related information available on mitigation behavior. Therefore, we provide the climate event probability and climate event occurrences as additional information in the observation space for both companies and investors. In the default 5-company-3-investor scenario, having additional climate-related information reduces the ending system climate risks, as shown in Figures 7(a). Figure 7(b) show similar effects of having additional climate-related information in the scenario where investors are not present. This result suggests that climate-related information helps both companies and investors make better environmental decisions. Even in the absence of ESG-focused investors, educating companies about the overall system improves outcomes. This aligns with literature showing that raising awareness encourages positive corporate responses (Delmas and Toffel, 2008; Bowen, 2000).
@@
In addition to the experiments presented above, we also explore settings where (1) company agents are allowed to spend on building their own resilience, (2) company agents are initialized with real-world company actions, (3) agents’ decisions are locked in for five years to simulate capital flexibility challenges, (4) companies’ climate resilience parameter $L_{t}^{\mathcal{C}_{i}}$ are random and vary across events and companies, and (5) a more strict bankruptcy mechanism. These experiments yield directional conclusions consistent with those presented in the main text. Detailed results can be found in Appendix 11.

### v3

@@ 4.1 The trade-off for the agents frames climate change as a social dilemma
Companies’ trade-off between short-term private costs and long-term collective benefits create a social dilemma within the environment. To illustrate this, we use empirical Schelling diagrams (Hughes et al., 2018), which plot the payoff of following either a cooperative or defecting policy, depending on the number of other cooperating company agents in a 5-company, 3-investor scenario.
@@
Figure 3: Schelling diagrams demonstrating that the environment constitutes a social dilemma. The graphs compare payoffs between cooperation (mitigation, blue lines) and defection (no mitigation, red lines) for a focal company, given varying number of other cooperating companies. Yellow lines represent the average payoff across all companies when the focal company defects. Subfigure (a) illustrates the selfish scenario, where all three investors consistently prioritize financial returns ($\alpha^{\mathcal{I}_{j}}=0,\text{ for }j=1,2,3$). Here, defection always yields higher payoffs for the focal company than cooperation, leading all companies to defect. However, widespread defection results in lower overall profits, as the average payoff (yellow) increases with greater cooperation, demonstrating the environment constitutes a social dilemma. Subfigure (b) and (c) correspond to two and three infinitely ESG-conscious investors ($\alpha^{\mathcal{I}_{j}}\approx\infty$), respectively. In (b), cooperation yields higher payoffs than defection for the focal company when few others cooperate. In (c), cooperation outperforms defection in all cases. Therefore, (b-c) demonstrate how investor behavior can transform the environment, eliminating the social dilemma by aligning corporate incentives with mitigation. Subfigures (d) and (e) build on (c) with three ESG-conscious investors. Subfigure (d) introduces resilience spending, while (e) adds greenwashing. The latter reintroduces a social dilemma, where corporations again avoid mitigation.
@@
In our case, cooperation represents altruistic investment in mitigating emissions. We simulate a cooperative policy as one which consistently invests 0.5% [^3] Since company actions are continuous, we selected 0.5% as a benchmark for illustrating the social dilemma structure in the Schelling diagram. This aligns with real examples of cooperative corporations, such as Patagonia, where $\sim$0.3of its valuation is allocated annually to climate initiatives. The diagram’s shape remains consistent across other tested values. of capital in mitigation and 0 in greenwashing or resilience building. In Figure 3(a) to 3(c), a defector policy takes zero action in all of mitigation, greenwashing, and resilience. As shown in Figure 3(a), when investors are solely profit-driven, the payoff for following a defector policy is always higher than the cooperator payoff. However, if all agents fail to cooperate, overall payoffs will be extremely low, showing that the environment represents a social dilemma. Figures 3(b) and 3(c) illustrate how the payoffs for cooperation and defection evolve as the proportion of ESG-conscious investors with positive ESG preferences increases. Notably, Figure 3(b) suggests the possibility of a bifurcated equilibrium when investor preferences are mixed. Specifically, climate-friendly companies may attract ESG-conscious investors, while other companies free ride and attract profit-motivated investors. Figure 3(c) shows that when all investors prioritize ESG-friendly firms, cooperation dominates defection in all cases. Therefore, with enough ESG-conscious investors the environment is no longer a social dilemma, and companies are actually greedily motivated to invest in mitigation to attract investors.
@@ 4.2 Multi-agent Reinforcement Learning Baselines
We are motivated to simulate companies and investors as independent, selfishly motivated agents that specialize in maximizing their own expected reward. We employ the state-of-the-art Independent PPO (IPPO) algorithm (De Witt et al., 2020; Yu et al., 2022) so each agent has its own policy parameters, and agents do not share parameters among themselves. InvestESG is implemented in both PyTorch (Paszke et al., 2017) and JAX (Bradbury et al., 2018), capable of simulating the investment behaviors in a massively parallel sense by vectorization on both CPUs and GPUs. By default, 5 companies are modeled with an initial capital of \$10 trillion each, alongside 3 investors, each starting with \$16 trillion. This creates a total beginning market wealth of \$98 trillion, which is comparable to the global stock market cap (WFE Statistics, 2023). We conduct the experiments in a 5-company, 3-investor environment by default unless specified otherwise. All the following results are averaged based on 3 runs of different random seeds, with the standard error plotted in the shaded region  [^5] In simulating climate events as described in Equation 2, we use a fixed random seed across training episodes. This reduces stochasticity while ensuring event occurrence varies with episode-specific climate risks. This design choice is based on two considerations. First, unlike MARL agents learning from scratch, real-world companies possess a baseline understanding of trade-offs and can reasonably anticipate future risks; a more predictable training environment mirrors this insight. Second, our study aims not to develop agents that solve climate change under all stochastic scenarios, but rather to efficiently train agents that grasp the strategic trade-offs faced by companies and investors. Limiting stochasticity makes learning and convergence easier without compromising the study’s structural conclusions. This design is also consistent with classic versions of climate-economic simulation models like DICE and RICE, where the link between economic development and $CO_{2}$ mass and the damage functions of global warming are largely deterministic (Nordhaus, 1992; Nordhaus & Yang, 1996; Nordhaus, 2017). The implementation also includes the option to not use a fixed random seed. . Following economic literature that argues more than 10 agents are often needed to observe meaningful group size effects (Arifovic et al., 2023), we provide additional results in Appendix 11 from settings with 10 companies and 10 investors, as well as 25 companies and 25 investors, which are consistent with our findings with the 5-company, 3-investor setting. [^6] The 25-by-25 agent setting aligns with Zhang et al. (2022), which employs 27 agents in a MARL framework to address climate change issues.
@@ 5 Main Results
Level of investors’ ESG-consciousness increases mitigation efforts. To evaluate the effectiveness of ESG disclosure mandates in incentivizing emissions mitigation by companies, we conduct three experiments: (1) Status Quo: All agents are profit-motivated, and no ESG scores are released; (2) Status Quo with Mandate: ESG scores are disclosed, but investors remain profit-driven ($\alpha=0$); (3) Mandate with ESG-Conscious Investors: ESG scores are disclosed, and investors are ESG-conscious ($\alpha>0$). In this first set of experiments, we study the default setting in which both greenwashing and resilience spending are disabled.
Figure 4: Ending values for all metrics averaged over the last 100 episodes; error bars show std. err. over 3 random seeds. We compare the status quo scenario with solely profit-driven investors (investors with ESG consciousness level of 0), both with and without the ESG disclosure mandate, to scenarios involving three ESG-conscious investors with ESG consciousness level of $\alpha=0.5$, $\alpha=1$, and $\alpha=10$. These results indicate that merely disclosing ESG scores is insufficient to resolve the social dilemma if investors are not interested in investing in climate-friendly companies. However as investors’ level of ESG consciousness increases, the ESG mandate results in consistent improvements in mitigation, climate risk, and market wealth.
The final system climate risks are shown in Figure 4. The values for Status Quo with Mandate and Status Quo are similar, indicating that mandatory ESG disclosure alone does not significantly incentivize mitigation when investors prioritize profits. However, we see that mitigation spending increases consistently corresponding to each increase in ESG-consciousness level ($\alpha=0.5,1,10$), leading to corresponding improvements climate risks and market wealth, with the best values achieved for $\alpha=10$. In this setting, a few leading companies attract the majority of ESG-conscious investments, forming a positive feedback loop where better market performance draws further investments (see Appendix Figure 11(c) and 11(d) for a demonstration for such pattern in the 10 companies, 10 investors case). These findings align with research showing that ESG disclosure mandates encourage climate-friendly efforts driven by societal and stakeholder pressures (Cormier & Magnan, 1999; Gamerschlag et al., 2011; Fama & French, 2007; Friedman & Heinle, 2016; Chen et al., 2018; Ioannou & Serafeim, 2019), and that fund managers are willing to sacrifice financial returns for ESG benefits (Krueger et al., 2021).
(a) Investment matrix at the latest episode for investors with different level of ESG consciousness: 0, 10, 10.
@@
Figure 5: Investigating the effects of the level of ESG consciousness in the case of 5 companies and 3 investors, where investor 0 is profit driven ($\alpha^{\mathcal{I}_{0}}=0$), and investors 1 and 2 are deeply climate-conscious ($\alpha^{\mathcal{I}_{1}}=\alpha^{\mathcal{I}_{2}}=10$). In (a) and (b), Company 0 (purple) learns to be the leading mitigator. The figures plot the investment distribution for each investor, showing that more climate-conscious investors focus on investing in the more climate-conscious companies, mirroring the market bifurcation results of the Schelling diagrams.
@@
Heterogeneous investor preferences lead to bifurcation in agent strategies. Research shows that investors have varying preferences for ESG efforts and respond in different ways (Amel-Zadeh & Serafeim, 2018). In the scenario plotted in Figure 5, we initialized three investors with different levels of ESG-consciousness, with Investor 0 representing the solely profit-seeking investor ($\alpha^{\mathcal{I}_{0}}=0$), and Investor 1 and 2 representing highly ESG-conscious investors ($\alpha^{\mathcal{I}_{1}}=\alpha^{\mathcal{I}_{2}}=10$), which corresponds to the scenario presented in the Schelling diagram in Figure 3(b). Figure 6(a) and 6(c) highlight a divergence in both investor and company behavior. Profit-driven Investor 0 distributes investments evenly across companies, while ESG-focused investors (1 and 2) favor climate-conscious firms, such as Company 0, which prioritizes mitigation. Such divergence extends to company strategies: Company 0 learns to attract more ESG investment by focusing on mitigation, while others prioritize financial returns at the expense of some investor interest.
(a) Mitigation and greenwashing amount for greenwashing coefficient = 2, 10, 20.
(b) Mitigation and greenwashing amount, zoomed in for early episodes.
@@
(c) Climate risk for various greenwashing coefficient, compared with status quo and status quo with esg=10.
@@
Figure 6: Mitigation, greenwashing, and climate risk for greenwashing coefficient = 2, 10, 20, with initial exploration episodes zoomed in in (b). It shows that when both companies and investors are IPPO agents, regardless of greenwashing cost, companies initially explore greenwashing and mitigation equally but quickly abandon greenwashing and invest in mitigation to attract ESG-conscious investors. In these experiments, the final climate risk is similar to the value when greenwashing is not enabled, suggesting greenwashing does not hinder mitigation investment.
@@
The possibility of greenwashing does not significantly undermine mitigation efforts. Building on research that examines the existence and extent of greenwashing (Wu et al., 2020; Marquis et al., 2016; Siano et al., 2017), and the concern that the ESG disclosure policy can backfire with greenwashing (El-Hage, 2021), we explore whether companies adjust their strategies when greenwashing is permitted. Based on the Schelling diagram in Figure 3(d) and Figure 6(c), we predict that investors would be misled by greenwashing, prompting companies to prioritize greenwashing over mitigation due to its lower cost, leading to waste of resources. To verify greenwashing can attract climate-conscious investors in our simulation, we hard-code companies to follow either a greenwashing or mitigation strategy. Figure 9 shows how in this scenario the greenwashing company earns a higher ESG score and succeeds in attracting climate-conscious investors. However, this trend is not observed when simulating agents with IPPO. Figure 6 plots results under varying greenwashing costs, represented by coefficient $\beta$ in Equation 3, where a larger $\beta$ indicates cheaper greenwashing. All investors have an ESG-consciousness level of $\alpha=1$. The results show that regardless of greenwashing cost, companies initially explore greenwashing and mitigation equally but quickly drop greenwashing as the main strategy. This likely occurs because, in the early episodes, investors have not yet linked ESG scores to their investment strategies, so do not respond to greenwashing efforts with increased investment. Instead, greenwashing presents an immediate cost for companies, which does not pay off with reduced climate risk, so they quickly learn to avoid it. In contrast, even without immediate investor rewards, companies may recognize the long-term benefits of mitigation (which leads to higher collective returns), and be incentivized to continue those efforts at first, even if they later learn to defect. This mirrors reality, where if investors are slow to adjust their strategies, companies may abandon greenwashing early but persist in low levels of mitigation for either its long-term climate benefits or the foresight of regulations and societal pressure that will eventually nudge them towards sustainable operation. This is consistent with empirical literature that despite the emergence of some greenwashing, ESG disclosure mandates overall encourage mitigation (Fiechter et al., 2022).
@@
Providing additional information about climate risk increases mitigation efforts. In this scenario, we test whether providing additional climate-related information to agents can affect mitigation behavior. Therefore, we provide the climate event probability and climate event occurrences as additional information in the observation space for both companies and investors. In the default 5-company-3-investor scenario, having additional climate-related information reduces the ending system climate risks, as shown in Figure 7(a). Figure 7(b) tests a scenario with no investors, in which companies are provided with the same additional climate-related information. Here climate information leads to a large reduction in climate risk, with the final risk ($\approx 0.7$) comparable to the risk achieved in the $\alpha=1$ scenario in Figure 4. This result suggests that climate-related information helps both companies and investors make better environmental decisions. Even in the absence of ESG-focused investors, educating companies about the overall system can significantly improve outcomes. This aligns with literature showing that raising awareness encourages positive corporate responses (Delmas & Toffel, 2008; Bowen, 2000).
@@
(a) Mitigation + resilience spending
(b) Climate risk
(c) Market total wealth
Figure 8: Effect of allowing resilience spending. When resilience is allowed, companies invest significantly more in resilience than in mitigation, resulting in ending episode higher market wealth, and lower climate risk.
Effects of resilience spending. In this case, we allow companies to invest in resilience spending to improve their own robustness to climate events, without affecting the global climate risk. Although investors continue to respond to companies’ ESG scores, resilience spending is excluded from the calculation of ESG scores. As shown in Figure 8, when resilience is allowed, companies invest significantly more in resilience than in mitigation. By investing in resilience, companies are more resilient to the climate event and therefore are able to maintain more capital to invest in actual mitigation, compared to the case when resilience is not allowed, reflected in 8(c). Therefore mitigation spending when resilience is allowed is actually slightly higher compared to when it is not an option, resulting in comparable final climate risk in Figure 8(b). This suggests that companies are financially incentivized to prioritize resilience investments, which can actually enable a greater commitment to climate mitigation efforts.
Additional experiments. In addition to the experiments presented above, we also explore settings where (1) company agents’ actions are initialized with data from real-world companies, (2) agents’ decisions are locked in for five years to simulate capital flexibility challenges, (3) companies’ climate resilience parameter $L_{t}^{\mathcal{C}_{i}}$ are random and vary across events and companies, and (4) the conditions for companies going bankrupt are made more strict. These experiments yield directional conclusions consistent with those presented in the main text. Detailed results can be found in Appendix 11.